OpenStudy (tiffanymak1996):

Solve X. A triangle ABC, with point D on line AB and point E on line AC. Point F is the intersection of line DC and EB. Angle ACD = 10 Angle DCB = 70 Angle EBC = 60 Angle EAD = 20 Angle EDC = x These r the only information given. Many Thanks!

OpenStudy (wolf1728):

Here's a graphic

imqwerty (imqwerty):



OpenStudy (sohailiftikhar):

yes wolfi done this already by telling the graph bro

OpenStudy (sohailiftikhar):

there is nothing to tell now

OpenStudy (wolf1728):

Angle ADC = 180 -20 -10 = 150 Degrees Angle CDB = 30 Degrees Angle B = 180 -20 -(10+70) = 80 Degrees Angle EBA = 20 Degrees Angle EFD = 360 -140 -150-20 = 50 Degrees = Angle CFB Well maybe I'll give this a rest now

OpenStudy (tiffanymak1996):

@imqwerty but x is Angle EDC, not Angle ADC

imqwerty (imqwerty):

oops :D sry i'll correct it

OpenStudy (tiffanymak1996):

thank you!

OpenStudy (tiffanymak1996):

@imqwerty i got everything you guys get, but i couldnt get EDC, that's why i'm asking it here. :P

OpenStudy (wolf1728):

Seems this is still unsolved. Let's see if this helps. Triangle ACB is isosceles side AC = side AB Triangle AEB is isosceles side AE = side EB Any ideas anyone?

OpenStudy (wolf1728):

Haven't given up yet. I redrew the graphic trying to be as accurate as possible. (All the angles were drawn as closely as possible to their posted values.)

OpenStudy (amilapsn):

\[\href{ https:///www.duckware.com/tech/worldshardesteasygeometryproblem.html }{\Huge \cal Click ~Here!!}\]

OpenStudy (wolf1728):

Thanks for the answer.

OpenStudy (amilapsn):

Yw!

OpenStudy (amilapsn):

But I recommend you not to read that answer straight away. Have sometime to try it yourself.

OpenStudy (wolf1728):

amilspan I've worked on it long enough. I drew both of the graphics posted here.

OpenStudy (tiffanymak1996):

thanks guys! Really appreciate the effort.:)))